OpenStudy (marlee):

2x/a=b+c solve for x

OpenStudy (anonymous):

multiply both sides by a and divide by 2 x=(ab+ac)/2

OpenStudy (anonymous):

x = [a(b+c)]/2

OpenStudy (saifoo.khan):

x = a(b+c) / 2

OpenStudy (llort):

x=0.5a(b+c) where a != 0